OpenStudy (anonymous):

Given the equation 4x + 2y = 9, which equation below would cause an inconsistent-independent system? A. 6x + 12y = 13 B. 2x + y = -6 C. x - 2y = 7 D. 8x - 4y = -5

OpenStudy (anonymous):

8x - 4y = -5 <- this one isn't it, the slope is different 2x + y = -6 (*2) 4x + 2y = -12 ^ this one looks exactly the same like the one in the problem statement it has a different solution even though it has the same amounts of y, x that means there's a contradiction and the system is inconsistent

OpenStudy (anonymous):

so which one would it be?

OpenStudy (anonymous):

B

OpenStudy (anonymous):

thnx :)

OpenStudy (anonymous):

yep